OpenStudy (anonymous):

How do I solve x + y = 3, x - y = 1 by graphing?

OpenStudy (anonymous):

first get y on its own for both equation

OpenStudy (anonymous):

so you would get y=-x+3 and y=x-1 agree?

OpenStudy (anonymous):

and then you would graph these two equation because they are in standard form y=mx+b m is the slope of line and b is y-intercept

OpenStudy (anonymous):

How do you get -x on the first one?

OpenStudy (anonymous):

x + y = 3 when u take x from left side to right u have to change the signs so since x is positive on left it would be negative on right